OpenStudy (anonymous):

Given the parent functions f(x) = log3 (5x − 5) and g(x) = log3 (x − 1), what is f(x) − g(x)? f(x) − g(x) = log3 (4x − 4) f(x) − g(x) = log3 (4x − 6) f(x) − g(x) = log3 5 f(x) − g(x) = = log3 one fifth

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <--- notice the 2nd rule listed there, so what do you think that gives us for \(\bf f(x) = log_3 (5x - 5) \qquad g(x) = log_3 (x - 1) \\ \quad \\ f(x)-g(x)=log_3 (5x - 5)-log_3 (x - 1)\implies ?\)

OpenStudy (anonymous):

would it be B f(x)-g(x)=log3(4x-6) @jdoe0001

OpenStudy (jdoe0001):

hmm what did you get for \(\bf log_3 (5x - 5)-log_3 (x - 1)\quad ?\)